OpenStudy (anonymous):

Factor the trinomial: 3x2 + 11x + 6 A. (3x + 2)(x + 3) B. (3x + 3)(x + 2) C. (3x + 1)(x + 6) D. (3x + 6)(x + 1)

OpenStudy (anonymous):

When you say "3x2" do you mean to say 3x^2?

OpenStudy (anonymous):

yes sorry @Pax

OpenStudy (anonymous):

Then I think the answer is (3x+2)(x+3)

OpenStudy (anonymous):

A

OpenStudy (anonymous):

I love factoring! Hold on let me solve this

OpenStudy (anonymous):

yeah A

OpenStudy (anonymous):

(3x + 2) • (x + 3)

OpenStudy (anonymous):

i'm so happy you love factoring @Aureyliant you are my new est friend haha

OpenStudy (anonymous):

2.1 Factoring 3x2+11x+6 The first term is, 3x2 its coefficient is 3 . The middle term is, +11x its coefficient is 11 . The last term, "the constant", is +6 Step-1 : Multiply the coefficient of the first term by the constant 3 • 6 = 18 Step-2 : Find two factors of 18 whose sum equals the coefficient of the middle term, which is 11 . -18 + -1 = -19 -9 + -2 = -11 -6 + -3 = -9 -3 + -6 = -9 -2 + -9 = -11 -1 + -18 = -19 1 + 18 = 19 2 + 9 = 11 That's it Step-3 : Rewrite the polynomial splitting the middle term using the two factors found in step 2 above, 2 and 9 3x2 + 2x + 9x + 6 Step-4 : Add up the first 2 terms, pulling out like factors : x • (3x+2) Add up the last 2 terms, pulling out common factors : 3 • (3x+2) Step-5 : Add up the four terms of step 4 : (x+3) • (3x+2) Which is the desired factorization Final result : (3x + 2) • (x + 3)

OpenStudy (anonymous):

Besties!! @cassiesilva0628

OpenStudy (anonymous):

Need anymore help? :)